OpenStudy (anonymous):

Help with trigonometry identities please! Simplify: a) cos(180º +x) + cos(180º - x) b) cos(7pi/126º)cos(36º/5)+sin(7pi/126º)sin(36º/5)

OpenStudy (amistre64):

cos(a+b) - cos(a-b): cos^2(a) - sin^2(a) + cos^2(a) + sin^2(a) 2cos^2(a) 2cos^2(180) ; cos(180) = -1 2(-1^2) = 2(1) = 2

OpenStudy (amistre64):

\[7\pi/126^O\] ????

OpenStudy (amistre64):

if anything that would equal 7pi(180)/126pi ... but ive never seen anything written like that before.

OpenStudy (anonymous):

this is the equation:

OpenStudy (amistre64):

well, its a cos(a-b) construct, but i dunno about those digits

OpenStudy (amistre64):

if you convert the degrees to radians, then I spoose you can have a radian measure without a pi in it.... but its just unnatural :)

OpenStudy (anonymous):

ok..thanks for the help on the first one

OpenStudy (amistre64):

7pi = 180 degrees, lets substitute that in instead

OpenStudy (amistre64):

cos(180/126)cos(36/5) + sin(180/126º)sin(36/5)

OpenStudy (amistre64):

then cos(180/126 - 36/5) is your angle....

OpenStudy (amistre64):

add the fractions and simplify... right?

OpenStudy (amistre64):

3636 ----- 630

OpenStudy (amistre64):

602/105...whats our remainder?

OpenStudy (anonymous):

did u already multiply the fraction by cos? im confused, sorry :S

OpenStudy (amistre64):

5 + (77/105)

OpenStudy (amistre64):

you dont need to do that :) you just need to find the angle of the cos: cos(a - b); find a - b

OpenStudy (amistre64):

108(5) 36(126) ------ - -------- 126(5) 126(5)

OpenStudy (amistre64):

180(5) 36(126) ------ - -------- 126(5) 126(5) arghh!!..typoed it lol

OpenStudy (amistre64):

3636 cos of ----- is what I get 630

OpenStudy (amistre64):

-3636 ------ you gotta catch this stuff, cause I dont always ;) 630

OpenStudy (anonymous):

Amistre please help :'(

OpenStudy (amistre64):

that looked alot better in my head lol

OpenStudy (amistre64):

loco..... give me a sec; im on break ;)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im not about to catch anything lol. i am very confused by trig in general..im about to post a question on trig identities...could u maybe help me through those when i post it? u seem like u know what ur doing :)

OpenStudy (amistre64):

its a farse, I no nutheeenng!!!

OpenStudy (anonymous):

he does. he know his stuff :)

OpenStudy (anonymous):

ur a superhero for a reason :)

OpenStudy (amistre64):

pfftt!!!.....super hero, now I gotta wear tights, and that aint a pretty picture ;)

OpenStudy (amistre64):

I get an angle of -5.77 degrees when we do all that.

OpenStudy (amistre64):

that is if its even right lol

OpenStudy (amistre64):

the cos of that angle = .9949 and thats the best I got fer ya on that one....right or wrong ;)

OpenStudy (anonymous):

r u solving the whole question btw? i just need to simplify..

OpenStudy (amistre64):

I solved it for the angle; andle = 5 + (77/105) and for the value of the cosine of that angle :=: .9949 how to simplify it is beyond me :)

OpenStudy (anonymous):

:'( ok. that's better than no answer at least!

OpenStudy (amistre64):

well, -5 and 77/105ths.... but yeah :) ..good luck with it; if anything, get it wrong and move ahead :)